OpenStudy (anonymous):

f(x)=x−6x^(1/3) x and y intercept of this function pls?

OpenStudy (anonymous):

Plug in x = 0 in the equation to get the y-intercept. Plug in y = 0 and solve for x to get the x-intercepts. Or do you need to answers to check your work?

OpenStudy (anonymous):

well i have tried the method of plugging in zeros and im not getting the right answer so i would like to see what people can come up with

OpenStudy (anonymous):

what can you come up with. i would like to ocmpare to what ive got

OpenStudy (mertsj):

If x = 0, y = 0 so the x and y intercepts are both 0

OpenStudy (anonymous):

the y-int is zero but the x-int isnt

OpenStudy (anonymous):

The y is straightforward. Is the y near x = 14?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

My solution for x is x = 6sqrt(6). It also corresponds with the graph of the function\[f(x) = x - 6x^\frac{1}{3}\]

OpenStudy (anonymous):

i came up with that ans earlier didnt work

OpenStudy (anonymous):

Factor out the function as:\[(x^\frac{2}{3} - 6)x^\frac{1}{3} = 0\]Both intercepts are x = 0 or x = 6sqrt6.

OpenStudy (anonymous):

i thought so too but i dont know why it doesnt work

OpenStudy (anonymous):

what about critical numbers?

OpenStudy (anonymous):

Is it an online answer? If so, have you tried other ways of writing the same number, like (6)^(3/2) or something like that? What I find weird is that the Wolf agrees that the intercepts are x = 0 and x = 6sqrt(6)

OpenStudy (anonymous):

i know...ive tried all those and none seem to work

OpenStudy (anonymous):

and its an online answer

OpenStudy (anonymous):

Very likely it's an error on the website or maybe bad formatting of the answer. But hey, someone could see another way of solving it. Anyway, good luck and I hope that you can solve the problem quickly :-)